# Analog In-memory Training on General Non-ideal Resistive Elements: The Impact of Response Functions

### v3

@@ Abstract
As the economic and environmental costs of training and deploying large vision or language models increase dramatically, analog in-memory computing (AIMC) emerges as a promising energy-efficient solution. However, the training perspective, especially its training dynamic, is underexplored. In AIMC hardware, the trainable weights are represented by the conductance of resistive elements and updated using consecutive electrical pulses. While the conductance changes by a constant in response to each pulse, in reality, the change is scaled by asymmetric and non-linear response functions, leading to a non-ideal training dynamic. This paper provides a theoretical foundation for gradient-based training on AIMC hardware with non-ideal response functions. We demonstrate that asymmetric response functions negatively impact Analog SGD by imposing an implicit penalty on the objective. To overcome the issue, we propose residual learning algorithm, which provably converges exactly to a critical point by solving a bilevel optimization problem. We demonstrate that the proposed method can be extended to address other hardware imperfections, such as limited response granularity. As we know, it is the first paper to investigate the impact of a class of generic non-ideal response functions. The conclusion is supported by simulations validating our theoretical insights.
@@ 1 Introduction
The remarkable success of large vision and language models is underpinned by advances in modern hardware accelerators, such as GPU, TPU [1], NPU [2], and NorthPole chip [3]. However, the computational demands of training these models are staggering. For instance, training LLaMA [4] cost \$2.4 million, while training GPT-3 [5] required \$4.6 million, highlighting the urgent need for more efficient computing hardware. Current mainstream hardware relies on the Von Neumann architecture, where the physical separation of memory and processing units creates a bottleneck due to frequent and costly data movement between them.
@@
Figure 1: The weight’s response curve. Positive and negative pulses are fired continuously on the left and right halves, respectively. One pulse is fired per cycle. Given $w$, the weight becomes $w^{+}$ or $w^{-}$ after one positive and negative pulse, respectively. The response factors $q_{+}(w)$ and $q_{-}(w)$ are approximately the slope of the curve at $w$, and $\Delta w_{\min}$ is the response granularity. (Left) Ideal response functions $q_{+}(w)\equiv q_{-}(w)$. Every point is a symmetric point. (Right) Asymmetric response functions $q_{+}(w)\neq q_{-}(w)$ almost everywhere expect for the symmetric point $w^{\diamond}$.
Despite its high efficiency, analog training is considerably more challenging than inference since it involves frequent weight updates. Unlike digital hardware, where the weight increment can be applied to the original weight in the memory cell, the weights in AIMC hardware are changed by the so-called pulse update. Pulse update. When receiving electrical pulses from its peripheral circuits, the resistive elements change their conductance in response to the pulse polarity [16]. Receiving a pulse at each pulse cycle, the conductance is updated by a small amount around $\Delta w_{\min}$, which is called response granularity. $\Delta w_{\min}$ is scaled by a response functions $q_{+}(w)$ or $q_{-}(w)$ depending on its polarity. Geometrically, $q_{+}(w)$ and $q_{-}(w)$ are the slopes of response curves; see Figure 1. All $\Delta w_{\min}$, $q_{+}(w)$, and $q_{-}(w)$ are element-specific parameters or functions that are set before training and hence remain fixed during training. Typically, $\Delta w_{\min}$ is known while $q_{+}(w)$ and $q_{-}(w)$ are not.
@@
where $\xi$ is a random data sample. Similar to stochastic gradient descent (SGD) in digital training (Digital SGD), the gradient-based training algorithm on AIMC hardware, Analog SGD, updates the weights by stochastic gradients $\nabla f(W_{k};\xi_{k})$. Digital SGD updates the weight by $W_{k+1}=W_{k}-\alpha\nabla f(W_{k};\xi_{k})$ with learning rate $\alpha$. Given a desired update $\Delta W=-\alpha\nabla f(W_{k};\xi_{k})$, AIMC hardware implements Analog SGD by sending about $|[\Delta W]_{i}|/\Delta w_{\min}$ pulses to the $i$-th element. With each pulse updating $W_{k}]_{i}$ by $\Delta w_{\min}$, $W_{k}]_{i}$ is ultimately updated by about $\Delta W]_{i}$.
Challenges of analog training. Despite its ultra-efficiency, gradient-based training on AIMC hardware is challenging. First, the generic response functions are asymmetric (i.e. $q_{+}(w)\not\equiv q_{-}(w)$), and non-linear [17, 18, 19]. Due to the variation of response functions and conductance states, gradients are scaled by different magnitudes (i.e., $q_{s}([W_{k}]_{i})$ for $s=\pm$) across different coordinates, leading to biased gradients. Furthermore, the response granularity $\Delta w_{\min}$ is a constant. When the gradients or the learning rate decay below $\Delta w_{\min}$, pulse update no longer provides sufficient precision to perform gradient descent [20]. Other imperfections include, but are not limited to, input/output (IO) of MVM operations and analog-digital conversion error [18]. This paper aims to investigate the impact of non-ideal response functions and develop a method to mitigate their negative effects. We also discuss extending the proposed method to deal with other hardware imperfections.
@@ 1.1 Main results
Recently, [21] partially answers the question by showing Analog SGD suffers from a convergence issue due to the asymmetric update, and a heuristic algorithm, Tiki-Taka [22, 23, 24], converges exactly by reducing the weight drift. However, their work is limited to a special case of linear response, which are in the form of $q_{+}(w)=1-w/{\tau},q_{-}(w)=1+w/{\tau}$ with hardware-specific parameter $\tau>0$. Given more general $q_{+}(w)$ and $q_{-}(w)$, the convergence of Tiki-Taka does not trivially holds, even though the response functions are still linear.
@@
Gap between theory for special linear and generic response. Consider a more generic linear response setting $q_{+}(w)=(1+c_{\texttt{Lin}})(1-w/{\tau}),q_{-}(w)=(1-c_{\texttt{Lin}})(1+w/{\tau})$ with a parameter $c_{\texttt{Lin}}$, which reduces to the setting in [21] when $c_{\texttt{Lin}}=0$. Figure 2 shows the damage from a non-zero $c_{\texttt{Lin}}$ to Tiki-Taka. Consistent with the conclusion in [21], Tiki-Taka significantly outperforms Analog SGD when $c_{\texttt{Lin}}=0$. However, when $c_{\texttt{Lin}}$ is perturbed from $0.1$ to $0.3$, Tiki-Taka degrades dramatically and even becomes worse than Analog SGD does. The modification is slight, but the convergence guarantee in [21] fails, and the convergence of Tiki-Taka is harmed significantly. This counter-example indicates a gap between the theory with special linear response and generic response, and necessitates the study of the analog training with generic response functions and the exploration of exact convergence conditions.
Ignoring other imperfections temporarily, we first analyze the impact of response functions. We show that Analog SGD suffers from asymptotic error due to the mismatch between the algorithmic stationary point and physical symmetric point. Inspired by that, we propose a novel algorithm framework that aligns two points, overcoming the asymmetric issues. Building on that, we endeavor to extend the proposed algorithm to more practical scenarios that involve other imperfections like limited granularity and noisy readings, prompting a second critical question:
@@ 1. C1)
    Building upon the equation of pulse update, we propose an approximated discrete-time dynamics of analog update. Enabled by it, we study the impact of response functions directly without being limited to concrete element candidates.
@@ 2. C2)
    Based on that, we show that instead of optimizing $f(\cdot)$, Analog SGD optimizes another penalized objective implicitly. An implicit penalty is introduced by the asymmetric response functions and attracts the weights towards symmetric points. Consequently, Analog SGD can only converge to the optimal point inexactly.
@@ 3. C3)
    We propose a novel Residual Learning algorithm framework to alleviate the asymmetric update and implicit penalty issues. Residual Learning explicitly introduces another residual array, which has a stationary point 0. By properly zero-shifting so that the stationary and symmetric points overlap, Residual Learning provably converges to a critical point.
@@ 4. C4)
    Building on C3), we propose a variant, Residual Learning v2, tailored for more practical training scenarios. We propose introducing a digital buffer to filter out the reading error due to the IO noise. Furthermore, we propose a threshold-based transfer rule to alleviate the instability due to the limited granularity.
@@ 2 Analog Training with Generic Response Functions
Compact formulations of analog update. We first investigate the dynamics of one element $w$ in $W\in{\mathbb{R}}^{D}$. This paper adopts $w$ to represent the element of the weight $W_{k}$ without specifying its index. The notation makes the formulations more concise and indicates that all elements are updated in parallel. As we discuss in Section˜1, the response granularity $\Delta w_{\min}$ is scaled by the response function $q_{s}(w)$. Naturally, given a desired update $\Delta w$, the increment is approximately scaled by $q_{s}(w)$ as well. Accordingly, we propose that an approximate dynamics of analog update is given by $w^{\prime}\approx U_{q}(w,\Delta w)$, where $U_{q}(w,\Delta w)$ is defined by
@@
The update (3) holds at each resistive element, while the model $W$ contains $D$ resistive elements with different response functions $q_{+}(\cdot)$ and $q_{-}(\cdot)$. We stack all the weights $w_{k}$ and expected increment $\Delta w_{k}$ together into vectors $W_{k},\Delta W_{k}\in{\mathbb{R}}^{D}$, where $k$ is the iteration index Similarly, the response functions $q_{+}(\cdot)$ and $q_{-}(\cdot)$ are stacked into $Q_{+}(\cdot)$ and $Q_{-}(\cdot)$, respectively. Let the notation $U_{Q}(W_{k},\Delta W)$ on matrices $W_{k}$ and $\Delta W$ denote the element-wise operation on $W_{k}$ and $\Delta W$, i.e. $U_{Q}(W_{k},\Delta W)]_{i}:=U_{[Q]_{i}}([W_{k}]_{i},[\Delta w]_{i}),\forall i\in[D]$ with $D]:=\{1,2,\cdots,D\}$ denoting the index set. The element-wise update (3) can be expressed as $W_{k+1}=U_{Q}(W_{k},\Delta W_{k})$. Leveraging the symmetric decomposition [22, 21], we decompose $Q_{-}(W)$ and $Q_{+}(W)$ into symmetric component $F(\cdot)$ and asymmetric component $G(\cdot)$
@@
We first introduce a commonly observed saturation property across a wide range of resistive elements. Resistive elements get saturated when they keep receiving the same pulses to avoid reaching arbitrarily high or low conductance states. Constrained by that, the conductance of the resistive element with $q_{+}(\cdot)$ and $q_{-}(\cdot)$ is bounded inside a dynamic range $\tau^{\min},\tau^{\max}]$ where $\tau^{\min}$ and $\tau^{\max}$ are the saturation points with zero responses, i.e. $q_{+}(\tau^{\max})=q_{-}(\tau^{\min})=0$. Near the saturation points, the asymmetric issue is significant, and thus, the update in one direction is suppressed, considerably impacting the convergence. On the contrary, if a point $w^{\diamond}$ satisfies $q_{-}(w^{\diamond})=q_{+}(w^{\diamond})$, the analog update behaves like a digital update. We refer to $w^{\diamond}$ as symmetric point. Symmetric points are typically located in the interior of the dynamic range and are far from saturation.
@@
Besides saturation, the function class should also enjoy other properties. First, the conductance ends up increasing when receiving a positive pulse and vice versa, leading to positive response functions. On top of that, we also assume the response functions are differentiable (and hence continuous) for mathematical tractability. Taking all into account, we define the following response function class.
@@
Definition˜1 covers a wide range of response functions, including but not limited to PCM, ReRAM, ECRAM, and others mention in Appendix˜A. Figure 3 showcases three examples from the response functions class, including linear, non-linear but monotonic, and even non-monotonic functions.
@@ 3 Implicit Penalty and Inexact Convergence of Analog SGD
This section introduces a critical impact of the response functions, implicit penalized objective. Affected by it Analog SGD can only converge inexactly with a non-diminishing asymptotic error.
@@ 3.1 Implicit penalty
Suppose ${\mathbb{E}}_{\xi_{k}}[|\nabla f(W_{k};\xi_{k})-{\mathbb{E}}_{\xi}[\nabla f(W_{k};\xi)]|]$ is a constant $\Sigma\in{\mathbb{R}}^{D}$ and let $D=1$. Analog SGD implicitly optimizes the following penalized objective
@@
The full version of Theorem 1 and its proof are deferred to Appendix G. In Theorem 1, $R_{c}(W)$ plays the role of a penalty to force the weight toward a symmetric point. As shown in Appendix G, $R_{c}(W)$ has a simple expression on linear response functions when $c_{\texttt{Lin}}=0$, leading (7) to $\min_{W}~f_{\Sigma}(W):=f(W)+\frac{\Sigma}{2\tau}\|W\|^{2}$ which is an $\ell_{2}$ regularized objective. In addition, the implicit penalty has a coefficient proportional to the noise level $\Sigma$ and inversely proportional to the dynamic range $\tau$. It implies that the implicit penalty becomes active only when gradients are noisy, and the noise amplifies the effect.

### v4

@@ Abstract
As the economic and environmental costs of training and deploying large vision or language models increase dramatically, analog in-memory computing (AIMC) emerges as a promising energy-efficient solution. However, the training perspective, especially its training dynamics, is underexplored. In AIMC hardware, the trainable weights are represented by the conductance of resistive elements and updated using consecutive electrical pulses. While the conductance changes by a constant in response to each pulse, in reality, the change is scaled by asymmetric and non-linear response functions, leading to a non-ideal training dynamics. This paper provides a theoretical foundation for gradient-based training on AIMC hardware with non-ideal response functions. We demonstrate that asymmetric response functions negatively impact Analog SGD by imposing an implicit penalty on the objective. To overcome the issue, we propose residual learning algorithm, which provably converges exactly to a critical point by solving a bilevel optimization problem. We demonstrate that the proposed method can be extended to address other hardware imperfections, such as limited response granularity. As we know, it is the first paper to investigate the impact of a class of generic non-ideal response functions. The conclusion is supported by simulations validating our theoretical insights.
@@ 1 Introduction
The remarkable success of large vision and language models is underpinned by advances in modern hardware accelerators, such as GPU, TPU [1], NPU [2], and NorthPole chip [3]. However, the computational demands of training these models are staggering. For instance, training LLaMA [4] cost \$2.4 million, while training GPT-3 [5] required \$4.6 million, highlighting the urgent need for more efficient computing hardware. Current mainstream hardware relies on the Von Neumann architecture, in which the physical separation of memory and processing units creates a bottleneck due to frequent, costly data movement between them.
@@
Figure 1: The weight’s response curve. Positive and negative pulses are fired continuously on the left and right halves, respectively. One pulse is fired per cycle. Given $w$, the weight becomes $w^{+}$ or $w^{-}$ after one positive and negative pulse, respectively. The response factors $q_{+}(w)$ and $q_{-}(w)$ are approximately the slope of the curve at $w$, and $\Delta w_{\min}$ is the response granularity. (Left) Ideal response functions $q_{+}(w)\equiv q_{-}(w)\equiv 1$. Every point is a symmetric point. (Right) Asymmetric response functions $q_{+}(w)\neq q_{-}(w)$ almost everywhere expect for the symmetric point $w^{\diamond}$.
Despite its high efficiency, analog training is considerably more challenging than inference since it involves frequent weight updates. Unlike digital hardware, where the weight increment can be applied to the original weight in the memory cell, the weights in AIMC hardware are changed by the so-called pulse update. Pulse update. When receiving electrical pulses from its peripheral circuits, the resistive elements change their conductance in response to the pulse polarity [16]. Receiving a pulse at each pulse cycle, the conductance is updated by $\Delta w_{\min}\cdot q_{+}(w)$ or $\Delta w_{\min}\cdot q_{-}(w)$, depending on the pulse polarity, where $\Delta w_{\min}$ is response granularity, and $q_{+}(w)$ and $q_{-}(w)$ are response functions. Geometrically, $q_{+}(w)$ and $q_{-}(w)$ are the slopes of response curves; see Figure 1. All $\Delta w_{\min}$, $q_{+}(w)$, and $q_{-}(w)$ are element-specific parameters or functions that are set before training and hence remain fixed during training. Typically, $\Delta w_{\min}$ is known while $q_{+}(w)$ and $q_{-}(w)$ are not.
@@
where $\xi$ is a random data sample. Similar to stochastic gradient descent (SGD) in digital training (Digital SGD), the gradient-based training algorithm on AIMC hardware, Analog SGD, updates the weights by stochastic gradients $\nabla f(W_{k};\xi_{k})$. Digital SGD updates the weight by $W_{k+1}=W_{k}-\alpha\nabla f(W_{k};\xi_{k})$ with learning rate $\alpha$. Given a desired update $\Delta W=-\alpha\nabla f(W_{k};\xi_{k})$, AIMC hardware implements Analog SGD by sending $|[\Delta W]_{d}|/\Delta w_{\min}$ pulses to the $d$-th element. Ideally, $q_{+}(w)=q_{-}(w)=1$ for every conductance states. If so, with each pulse updating $W_{k}]_{d}$ by $\Delta w_{\min}$, $W_{k}]_{d}$ is ultimately updated by about $\Delta W]_{d}$.
Challenges of analog training. Despite its ultra-efficiency, gradient-based training on AIMC hardware is challenging. First, the generic response functions are asymmetric (i.e. $q_{+}(w)\not\equiv q_{-}(w)$), and non-linear [17, 18, 19]. Due to the variation of response functions and conductance states, gradients are scaled by different magnitudes across different coordinates, leading to biased gradients. Furthermore, the response granularity $\Delta w_{\min}$ is a constant. When the gradients or the learning rate decay below $\Delta w_{\min}$, pulse update no longer provides sufficient precision to perform gradient descent [20]. Other imperfections include, but are not limited to, noisy input/output (IO) of MVM operations and analog-digital conversion error [18]. This paper aims to investigate the impact of non-ideal response functions and develop a method to mitigate their negative effects. We also discuss extending the proposed method to deal with other hardware imperfections.
@@ 1.1 Main results
Recently, [21] partially answers the question by showing that Analog SGD suffers from a convergence issue due to the asymmetric update, and a heuristic algorithm, Tiki-Taka [22, 23, 24], converges exactly by reducing the weight drift. However, their work is limited to a special case of linear response functions, which are in the form of $q_{+}(w)=1-w/{\tau},q_{-}(w)=1+w/{\tau}$ with hardware-specific parameter $\tau>0$. Given more general $q_{+}(w)$ and $q_{-}(w)$, the convergence of Tiki-Taka does not trivially hold, even though the response functions are still linear.
@@
Gap between theory for special linear and generic response functions. Consider a more generic linear response $q_{+}(w)=(1+c_{\texttt{Lin}})(1-w/{\tau}),q_{-}(w)=(1-c_{\texttt{Lin}})(1+w/{\tau})$ with a parameter $c_{\texttt{Lin}}$, which reduces to the setting in [21] when $c_{\texttt{Lin}}=0$. Figure 2 shows the damage from a non-zero $c_{\texttt{Lin}}$ to Tiki-Taka. Consistent with the conclusion in [21], Tiki-Taka significantly outperforms Analog SGD when $c_{\texttt{Lin}}=0$. However, when $c_{\texttt{Lin}}$ is perturbed from $0.1$ to $0.3$, Tiki-Taka degrades dramatically and even becomes worse than Analog SGD does. The modification is slight, but the convergence guarantee in [21] fails, and the convergence of Tiki-Taka is harmed significantly. This counter-example indicates a gap between the theory for special linear and generic response functions, and necessitates the study of the analog training with generic response functions and the exploration of exact convergence conditions.
Ignoring other imperfections temporarily, this paper first analyzes the impact of response functions. We show that Analog SGD suffers from asymptotic error due to the mismatch between the algorithmic stationary point and physical symmetric point. Inspired by that, we propose a novel algorithm framework that aligns two points, overcoming the asymmetric issues. Building on that, we endeavor to extend the proposed algorithm to more practical scenarios that involve other imperfections like limited granularity and noisy readings, prompting a second critical question:
@@ 1. C1)
    Building on the pulse update equation, we propose an approximate discrete-time dynamics for analog update. Enabled by this, we study the impact of response functions directly, without being limited to specific element candidates.
@@ 2. C2)
    Based on that, we show that instead of optimizing $f(\,\cdot\,)$, Analog SGD optimizes another penalized objective implicitly. An implicit penalty is introduced by the asymmetric response functions, which attract the weights towards symmetric points. Consequently, Analog SGD can only converge to the optimal point inexactly.
@@ 3. C3)
    We propose a novel Residual Learning theoretical framework to alleviate the asymmetric update and implicit penalty issues. Residual Learning explicitly introduces another residual array, which has a stationary point 0. This framework leads to Tiki-Taka heuristically proposed in [22] while it offers an understanding of how Tiki-Taka deals with the challenge from generic response functions. By properly zero-shifting so that the stationary and symmetric points overlap, Residual Learning provably converges to a critical point.
@@ 4. C4)
    Building on C3), we propose a variant, Residual Learning v2, tailored for more practical training scenarios. We propose introducing a digital buffer to filter out reading errors caused by IO noise. Furthermore, we propose a threshold-based transfer rule to alleviate instability caused by limited granularity.
@@ 2 Analog Training with Generic Response Functions
Compact formulations of analog update. We first investigate the dynamics of one element $w$ in $W\in{\mathbb{R}}^{D}$. This paper adopts $w$ to represent the element of the weight $W_{k}$ without specifying its index. As we discuss in Section˜1, the response granularity $\Delta w_{\min}$ is scaled by the response functions $q_{+}(w)$ or $q_{-}(w)$. Since a desired update $\Delta w$ requires a series of pulses with each scaled by approximately $q_{+}(w)$ or $q_{-}(w)$, it is sensible that the $\Delta w$ is approximately scaled by $q_{+}(w)$ or $q_{-}(w)$ as well. Accordingly, we propose that an approximate dynamics of analog update is given by $w^{\prime}\approx U_{q}(w,\Delta w)$, where $U_{q}(w,\Delta w)$ is defined by
@@
The update (3) holds at each resistive element. At the $k$-th iteration, We stack all the weights $w_{k}$ and expected increment $\Delta w_{k}$ together into vectors $W_{k},\Delta W_{k}\in{\mathbb{R}}^{D}$. Similarly, the response functions $q_{+}(\,\cdot\,)$ and $q_{-}(\,\cdot\,)$ are stacked into $Q_{+}(\,\cdot\,)$ and $Q_{-}(\,\cdot\,)$, respectively. Let the notation $U_{Q}(W_{k},\Delta W)$ on matrices $W_{k}$ and $\Delta W$ denote the element-wise operation on $W_{k}$ and $\Delta W$, i.e. $U_{Q}(W_{k},\Delta W)]_{d}:=U_{[Q]_{d}}([W_{k}]_{d},[\Delta w]_{d}),\forall d\in[D]$ with $D]:=\{1,2,\cdots,D\}$ denoting the index set. The element-wise update (3) can be expressed as $W_{k+1}=U_{Q}(W_{k},\Delta W_{k})$. Leveraging the symmetric decomposition [21, 22], we decompose $Q_{-}(W)$ and $Q_{+}(W)$ into symmetric component $F(\,\cdot\,)$ and asymmetric component $G(\,\cdot\,)$
@@
We first introduce a commonly observed saturation property across a wide range of resistive elements. Resistive elements get saturated when they keep receiving the same pulses to avoid reaching arbitrarily high or low conductance states. Constrained by that, the conductance of the resistive element with $q_{+}(\,\cdot\,)$ and $q_{-}(\,\cdot\,)$ is bounded inside a dynamic range $\tau^{\min},\tau^{\max}]$ where $\tau^{\min}$ and $\tau^{\max}$ are the saturation points with zero responses, i.e. $q_{+}(\tau^{\max})=q_{-}(\tau^{\min})=0$. Near the saturation points, the asymmetric issue is significant, so the update in one direction is suppressed, significantly impacting convergence. On the contrary, if a point $w^{\diamond}$ satisfies $q_{-}(w^{\diamond})=q_{+}(w^{\diamond})$, the analog update behaves like a digital update. We refer to $w^{\diamond}$ as symmetric point. Symmetric points are typically located in the interior of the dynamic range and are far from saturation.
@@
Besides saturation, the function class should also enjoy other properties. First, the conductance increases when receiving a positive pulse and vice versa, leading to positive response functions. On top of that, we also assume the response functions are differentiable (and hence continuous) for mathematical tractability. Taking all into account, we define the following response function class.
@@
Definition˜1 covers a wide range of response functions, including but not limited to PCM, ReRAM, ECRAM, and others mentioned in Appendix˜A. Figure 3 showcases three examples from the response functions class, including linear, non-linear but monotonic, and even non-monotonic functions.
@@ 3 Implicit Penalty and Inexact Convergence of Analog SGD
This section introduces a critical impact of the response functions, implicit penalized objective. Affected by this, Analog SGD can only converge inexactly with a non-diminishing asymptotic error.
@@ 3.1 Implicit penalty
Assume ${\mathbb{E}}_{\xi_{k}}[|\nabla f(W_{k};\xi_{k})-{\mathbb{E}}_{\xi}[\nabla f(W_{k};\xi)]|]$ is a constant $\Sigma$ and suppose that $D=1$. Analog SGD implicitly optimizes the following penalized objective
@@
The full version of Theorem 1 and its proof are deferred to Appendix G. In Theorem 1, $R_{c}(W)$ plays the role of a penalty to force the weight towards a symmetric point. As shown in Appendix G, $R_{c}(W)$ has a simple expression on linear response functions when $c_{\texttt{Lin}}=0$, leading (7) to $\min_{W}~f_{\Sigma}(W):=f(W)+\frac{\Sigma}{2\tau}\|W\|^{2}$ which is an $\ell_{2}$ regularized objective. In addition, the implicit penalty has a coefficient proportional to the noise level $\Sigma$ and inversely proportional to the dynamic range $\tau$. It implies that the implicit penalty becomes active only when gradients are noisy, and the noise amplifies the effect.